Circular No. 8278

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V445 PUPPIS D. K. Lynch, R. J. Rudy, S. Mazuk, and C. C. Venturini, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.5–2.5-μm spectroscopy of V445 Pup on Jan. 14 and 16 UT using the Aerospace Corporation’s Near-Infrared Imaging Spectrograph on the Lick 3-m telescope. The object has faded to fainter than mag J = 18 and could not be detected spectroscopically in the visible. In the infrared, only the He-I 1.0830- and 2.0581-μm lines were seen, both showing doubled profiles, as though a bipolar outflow is taking place. Notably absent were He II or coronal lines, suggesting that the ionizing source is not hot enough to produce significant He II or the more highly ionized states associated with the coronal features. The very red continuum detected longward of 1.5 μm appears to be produced by emission from hot dust and affected by reddening. Additional magnitudes: H = 16.7 ± 0.2, K_{s} [2.16-μm] = 13.8 ± 0.2. In view of this very unusual behavior, further monitoring is recommended.

COMETS C/2003 B2, 2003 C2, 2003 C3, 2003 C4, AND 2003 C5 (SOHO)

Further to IAUC 8276, K. Battams reports measures of additional Kreutz comets found on SOHO C3 website images by D. Evans (C/2003 C2, which was also visible in C2-coronagraph images), R. Kracht (C/2003 C3), and J. Sachs (C/2003 C4 and C/2003 C5). Reduced positions and orbital elements by B. G. Marsden appear on MPEC 2004-B34.

Corrigendum. On IAUC 8276, the discovery date for C/2003 B2 should read Jan. 18.821.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 C2 Feb. 8.433  21^{h}57.^{m}3  −17^{o}29′
C/2003 C3  10.613  21 51.5  −15 30
C/2003 C4  13.929  22 05.6  −14 16
C/2003 C5  15.971  22 21.0  −14 19

SUPERNOVA 2004B IN IC 390 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004B (cf. IAUC 8268), obtained by P. Berlind on Jan. 28.16 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks past maximum.

2004 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green